Your bank will lend you $3,600 for 55 days at a cost of $96 interest.  
  
What is your effective rate of interest? (Use a 360-day year. Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
Answer #1

Int Rate per 55 days = Int / Principal

= $ 96 / $ 3600

= 0.0267 i.e 2.67%

EAR = (1+r)^n - 1

n = 360 / 55

= 6.55

r is int rate per 55 days

EAR = (1+r)^n - 1

= ( 1 + 0.0267 )^ 6.55 - 1

= ( 1.0267 ^ 6.55) - 1

= 1.1880 - 1

= 0.1880 i.e 18.80%

Effective Rate of Int is 18.80%
Answer #2

SOLUTION :



Effective annual rate

= [ ( 1 + interest rate of the period)^ (number of periods in a year) ] - 1

= (1 + 96/3600)^(360/55) - 1

= 0.1880

= 18.80 % (ANSWER).


answered by: Tulsiram Garg